OpenStudy (anonymous):

help me plzz @AriPotta @AMYCARTER @Hotchellerae21 @mad

OpenStudy (anonymous):

QUESTION?

OpenStudy (anonymous):

Match each shape with its surface area. Use three point one four for pi and round to the nearest whole number where necessary. Match Term Definition square pyramid: base length 6 meters, slant height ten meters A) one hundred fifty-one square meters cone: radius 4 meters, slant height 8 meters B) one hundred fifty-six square meters square pyramid: base length 5 meters, slant height eleven meters C) one hundred eighty-eight square meters cone: radius 6 meters, slant height 4 meters D) one hundred thirty-five square meters

OpenStudy (anonymous):

square pyramid: SA = B + (1/2)pL = (10 m)^2 + (1/2)(40 m)(6 m) = 100 + 120 = 220 m^2 B cone: SA = πr^2 + πrL ≈ (3.14)(8 m)^2 + (3.14)(8 m)(4 m) = 3.14(64 + 32) = 301.44 m^2 C square pyramid: SA = (11 m)^2 + (1/2)(44 m)(5 m) = 121 + 110 = 231 m^2 D cone: SA = (3.14)(5 m)^2 + (3.14)(5 m)(7 m) = 3.14(25 + 35) = 188.4 m^2 A